# A way around the exponential scaling in optimal quantum control

### v1

@@ Appendix B Numerical optimisation
The evolution of $\bm{a}(t)$ is obtained with a Krylov subspace method to take advantage of the sparsity of $K(t)$. We feed the infidelity and its gradient with respect to the control variables to a gradient-based optimisation algorithm. The interior point method is employed, and the Hessian is approximated with the L-BFGS method. This algorithm is implemented in the fmincon function in Matlab.
@@
For large $n$, there are many shallow local minima and plateaus in the control landscape, especially for $U_{G}$ due to the difficulty of reaching the $n$-body interaction term. To systematically address these challenges, a three-step procedure is employed. Initially, the symmetry of the control problem is imposed on the control variables, followed by the relaxation of this symmetry. Specifically, we begin with
@@
We then lower the infidelity for each $n$ by optimising with a control Hamiltonian with inversion symmetry
@@
$+h(t)(X_{1}+X_{n}),$ (56)
using the final control in the previous step as the initial guess. Finally, we refine the solution further by optimizing with a control Hamiltonian that lacks any symmetry
$H^{c}_{3}(t)=g\!\sum_{j=1}^{n-1}X_{j}X_{j+1}\!+\!\sum_{j=1}^{n}f_{j}(t)Z_{j}\!+\!\!\!\!\sum_{j\in\{1,n\}}\!\!\!\!h_{j}(t)X_{j}.$ (57)
@@
For the optimisations of $H_{C},H_{D},H_{G}$, $T\approx n\pi/2g$ and the number of time bins in the piece-wise control pulse is $M=10n$. For $U_{C}$ and $U_{D}$, $T\approx\pi/2g$ and $M=10n$. For $U_{G}$, $T\approx 2n\pi/g$ and $M=20n$.
@@
Appendix C Lie algebra of the hexagonal spin ladder

### v2

@@ B.3 Optimisation
We feed the infidelity and its gradient with respect to the control variables to a gradient-based optimisation algorithm. The interior point method is employed, and the Hessian is approximated with the L-BFGS method. This algorithm is implemented in the fmincon function in Matlab.
@@
For large $n$, there are many shallow local minima and plateaus in the control landscape for $U_{G}$ due to the difficulty of reaching the $n$-body interaction term. To systematically address these challenges, a three-step procedure is employed. Initially, the symmetry of the control problem is imposed on the control variables, followed by the relaxation of this symmetry. Specifically, we begin with
@@
We then lower the infidelity for each $n$ by optimising with a system Hamiltonian with inversion symmetry
@@
$+w(t)(X_{1}+X_{n})\ ,$ (68)
using the final control in the previous step as the initial guess. Finally, we refine the solution further by optimizing with a system Hamiltonian that lacks any symmetry
$H^{c}_{3}(t)=g\!\sum_{j=1}^{n-1}X_{j}X_{j+1}\!+\!\sum_{j=1}^{n}f_{j}(t)Z_{j}\!+\!\!\!\!\sum_{j\in\{1,n\}}\!\!\!\!w_{j}(t)X_{j}\ .$ (69)
@@
$T\approx n\pi/2g$ for the optimisations of $H_{G},H_{C}$, $T\approx n\pi/g$ for $H_{D}$, and the number of time bins in the piece-wise control pulse is $M=10n$ for all three cases. For $U_{C}$ and $U_{D}$, $T\approx 2\pi g$ and $M=10n$. For $U_{G}$, $T\approx 2n\pi/g$ and $M=20n$.
B.4 Adiabatic error estimate
The control target $\mathcal{I}_{D}$ for $H_{D}$ is defined as the operator $\mathcal{I}(T_{a})$ resulting from simulated adiabatic evolution with a very long duration ($T_{a}=2500\tau_{g}$) from the initial operator $\mathcal{I}(0)$, as described in Sec. B.1.1. In a perfect adiabatic evolution, each eigenstate of ${\cal I}(0)$ would evolve to a corresponding eigenstate of $H_{T}$ and, thus, ${\cal I}(T_{a})$ would commute with $H_{T}$. To verify this commutativity, one can propagate ${\cal I}(t)$ with $H_{T}$ for an additional time $T_{b}$ randomly chosen from the interval $[50\tau_{g},100\tau_{g}]$. In the ideal adiabatic limit, since ${\cal I}(T_{a})$ commutes with $H_{T}$, further evolution would not change the system, resulting in ${\cal I}(T_{a}+T_{b})={\cal I}(T_{a})$. Any deviation from this equality is a measure of adiabatic error, and the infidelity $1-\mbox{tr}({\cal I}(T_{a}){\cal I}(T_{a}+T_{b}))/\mbox{tr}({\cal I}(T_{a})^{2})$ can be used as an estimate of this error.
B.5 MPS verification
@@
Appendix C Relation between the state infidelity bound and the operator infidelity
Given the system Hamiltonian $H(t)$ (Eq. (20)), the initial condition ${\cal I}(0)=\sum_{j}Z_{j}$ and a control target ${\cal I}_{T}$ that coincides with the target Hamiltonian $H_{T}$ (which includes the cases of $H_{G}$ and $H_{C}$ defined in Sec. III.1), the bound
$\mathcal{B}=\frac{\bra{\Psi(T)}H_{T}\ket{\Psi(T)}-E_{0}}{E_{1}-E_{0}}$ (70)
in the state infidelity defined in Eq.  (18) depends on the operator infidelity $\mathcal{J}$ via the linear relation
$\mathcal{B}=\frac{n\mathcal{J}-s}{E_{1}-E_{0}}\ ,$ (71)
with a term $s$ that, under assumptions on typicality, becomes vanishingly small in the limit $n\to\infty$.
This is proven in the following.
Given the requirement that ${\cal I}_{T}=H_{T}$, the expectation value in Eq. (70) satisfies the equality
$\bra{\Psi(T)}H_{T}\ket{\Psi(T)}=\bra{\Psi(0)}\tilde{\mathcal{I}}(T)\ket{\Psi(0)}\ ,$ (72)
where ~$\tilde{\mathcal{I}}(T)=U^{\dagger}(T)\mathcal{I}_{T}U(T)$, |$\ket{\Psi(0)}$ is the ground state of $\mathcal{I}(0)$, and $E_{0}$ the ground energy of both $H_{T}$ and $\mathcal{I}(0)$. Using the relations
$\mbox{tr}(\mathcal{I}(T)\mathcal{I}_{T})=\mbox{tr}(U(T)\mathcal{I}(0)U^{\dagger}(T)\mathcal{I}_{T})=\mbox{tr}(\mathcal{I}(0)\tilde{\mathcal{I}}(T))$ (73)
and $\mbox{tr}(\mathcal{I}_{T}^{2})=\mbox{tr}(\mathcal{I}^{2}(0))$, the operator infidelity in Eq. (12) can be expressed as
$\mathcal{J}=1-\frac{\mbox{tr}(\mathcal{I}(0)\tilde{\mathcal{I}}(T))}{\mbox{tr}(\mathcal{I}^{2}(0))}\ .$ (74)
If the control target is reached exactly, the identity ~$\tilde{\mathcal{I}}(T)=\mathcal{I}(0)$ holds, and any imperfection of the optimal control solution can be captured by the deviation
$\Delta\mathcal{I}=\mathcal{I}(0)-\tilde{\mathcal{I}}(T)\ .$ (75)
The deviation $\Delta\mathcal{I}$ can be expanded in terms of the operators $\mathfrak{a}_{j}$ (Eqs. (21j)) as $\Delta\mathcal{I}=\sum_{j}c(\mathfrak{a}_{j})\mathfrak{a}_{j}$, with expansion coefficients
[…]
The operator infidelity can be expressed as
$\mathcal{J}=\frac{\mbox{tr}(\mathcal{I}(0)\Delta\mathcal{I})}{\mbox{tr}(\mathcal{I}^{2}(0))}\ .$ (77)
With $\mathcal{I}(0)=\sum_{j}Z_{j}$, $\mbox{tr}(\mathcal{I}^{2}(0))=n2^{n}$, and the overlaps $\mbox{tr}(\mathcal{I}(0)Z_{j})=2^{n}$, and $\mbox{tr}(\mathcal{I}(0)\mathfrak{a}_{j})=0$ for all operators $\mathfrak{a}_{j}$ that are not single-spin $Z$ operators (i.e. Eqs. (21b) to (21j)), the operator infidelity reduces to
$\mathcal{J}=\frac{\sum_{j}c(Z_{j})}{n}\ ,$ (78)
in terms of the expansion coefficients $c(Z_{j})$ defined in Eq. (76).
Given the ground state energy $\bra{\Psi(0)}\mathcal{I}(0)\ket{\Psi(0)}=E_{0}$, the bound ${\cal B}$ can be expressed as
$\mathcal{B}=-\frac{\bra{\Psi(0)}\Delta\mathcal{I}\ket{\Psi(0)}}{E_{1}-E_{0}}\ .$ (79)
With $\Delta\mathcal{I}$ expanded in terms of the operators $\mathfrak{a}_{j}$ (Eqs. ((21j)) and the expectation values $\bra{\Psi(0)}Z_{j}\ket{\Psi(0)}=-1$, $\bra{\Psi(0)}\prod_{j}Z_{j}\ket{\Psi(0)}=(-1)^{n}$, and $\bra{\Psi(0)}\mathfrak{a}_{j}\ket{\Psi(0)}=0$ for all other cases (i.e. Eqs. (21b) to (21i)), this yields
$\mathcal{B}=\frac{\sum_{j}c(Z_{j})-s}{E_{1}-E_{0}}\ ,$ (80)
where $s=(-1)^{n}c(\prod_{j}Z_{j})$, and the expansion coefficient $c(\prod_{j}Z_{j})$ is defined in Eq. (76). Together with Eq. (78), this results in Eq. (71).
While it is conceivable that $\Delta{\cal I}$ is proportional to the operator $\prod_{j}Z_{j}$, one would expect that, typically, $\Delta{\cal I}$ would have contributions from all operators $\mathfrak{a}_{j}$ in Eq. (21j). In estimating the magnitude of $s$, we will thus assume the expansion coefficients $c(\mathfrak{a}_{j})$ to be random with each coefficient resulting from a distribution with zero mean and standard deviation $\delta$. The typical magnitude of the term
$n\mathcal{J}=\sum_{j=1}^{n}c(Z_{j})$ (81)
is then given by its standard deviation, i.e. by $\sqrt{n}\delta$. The typical magnitude of $c\left(\prod_{j}Z_{j}\right)$, on the other hand, is given by $\delta$. The term $s$ thus becomes negligible as compared to the term $n\mathcal{J}$ for large systems.
Fig. 7 depicts the deviation of the actual results of Fig. 2a from the linear dependence $\mathcal{B}=\frac{n\mathcal{J}}{E_{1}-E_{0}}$ as a function of the system size $n$. The figure shows that the deviation decays rapidly with $n$. The actual results thus follow the linear dependence in good approximation.
Figure 7: Deviation $D=\left|1-(E_{1}-E_{0})\mathcal{B}/n\mathcal{J}\right|=|s|/n\mathcal{J}$, in log-scale, from the linear relation between the bound $\mathcal{B}$ on state infidelity and operator infidelity $\mathcal{J}$ that is expected to hold in the limit $n\to\infty$ following Eq. (71). The deviation obtained from the data shown in Fig. 2a decays rapidly with $n$, consistently with Eq. (71).
Appendix D Lie algebra of the hexagonal spin ladder